OpenStudy (anonymous):

solve e^2x-3e^x+2=0 for x

OpenStudy (callisto):

\[e^{2x}-3e^x+2=0\]\[(e^x)^2-3e^x+2=0\]Quadratic equation. Let u = e^x, you'll get u^2 - 3u +2 =0 Solve u, then solve x

OpenStudy (anonymous):

where did u come from tho?

OpenStudy (anonymous):

where did the t come from as well?

OpenStudy (callisto):

You let u (or t) be e^x, just easier for you to see it is a quadratic equation. Of course, you don't have to.

OpenStudy (anonymous):

is there a simplier way to solve it?

OpenStudy (callisto):

This is very simple.

OpenStudy (callisto):

Using factorization: \[e^{2x}-3e^x+2=0\]\[(e^{x}-2)(e^x-1)=0\]Almost done.

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

do i find the zeros?

OpenStudy (callisto):

Hmm.. Solve for e^x.

OpenStudy (anonymous):

well see, im not too sure how to solve for that

OpenStudy (callisto):

by putting each factor =0

OpenStudy (anonymous):

is that where e equals to something?

OpenStudy (callisto):

e^x equal to something. Then, take natural log for both sides.

OpenStudy (anonymous):

Actually, you are solving for x in e^x. e is a constant, so you arr trying to find the value x so that when e is raised to the power, you get the value 1 or 2. In other words, \[e^{x}=2 \] or \[e^{x}=1\]

OpenStudy (anonymous):

And then just do as Callisto suggested, take the natural log of both sides.

OpenStudy (anonymous):

oh okay, im a bit confused but i think i understand

OpenStudy (anonymous):

Where is the confusion?

OpenStudy (anonymous):

why is raised to 1 or 2? @calmat01